OpenStudy (anonymous):

whats the numerator for the following rational expression? a/b + 8/b = ?/b

OpenStudy (anonymous):

a+8

mathslover (mathslover):

\[\huge{\frac{a}{b}+\frac{8}{b}=\frac{a+8}{b}}\] hence numerator = \(\huge{a+8}\)